# Mapping Cone Connections and their Yang-Mills Functional

### v1

@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
For each $\gamma\in\Psi^{a}_{a,\zeta}(M)$, there exists some disk $D$ such that $\partial D=\gamma$, and $\int_{D}\zeta=0$. By Lemma 4.5, the holonomy along $\gamma$ is $hol(\gamma)=-\mathrm{exp}\Big{(}\int_{D}\zeta\cdot\xi\Big{)}=e$. So we can define a morphism
$\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G,\quad\gamma\mapsto hol(\gamma).$
We will show that $(P_{\rho},A_{\rho})$ and $(P,A)$ are equivalent.
Given $[\delta,g]\in P_{\rho}$, let $\tilde{\delta}$ be the horizontal lift of $\delta$ in $P$ with $\tilde{\delta}(0)=u_{0}$. Set
$f:P_{\rho}\to P,\quad[\delta,g]\mapsto\tilde{\delta}(1)g.$
Suppose $[\delta_{1},g_{1}]$ and $[\delta_{2},g_{2}]$ represent the same class in $P_{\rho}$, then there exists $\gamma_{0}\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ such that $\delta_{1}=\delta_{2}\gamma_{0}$ in $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ and $g_{1}=\big{(}\rho(\gamma_{0})\big{)}^{-1}g_{2}$ in $G$. So $\gamma_{0}^{-1}\delta_{2}^{-1}\delta_{1}\in\Psi^{a}_{a,\zeta}(M)$, and its holonomy is trivial according to Lemma 4.5. Thus, the holonomy of $\delta_{2}^{-1}\delta_{1}$ is $hol(\gamma_{0})=\rho(\gamma_{0})$. Let $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ denote the horizontal lift of $\delta_{1}$ and $\delta_{2}$, respectively, starting at $u_{0}$. Since $\tilde{\delta}_{1}$ and the horizontal lift of $\delta_{2}\gamma_{0}=\delta_{2}\delta_{2}^{-1}\delta_{1}$ starting at $u_{0}$ have the same ending point, so do $\tilde{\delta}_{1}$ and the horizontal lift of $\delta_{2}$ starting at $u_{0}\cdot hol(\gamma_{0})$. Then the ending points of $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ satisfy $\tilde{\delta}_{1}(1)=\tilde{\delta}_{2}(1)\cdot hol(\gamma_{0})$. Therefore, by definition,
@@
Hence, $f$ is well defined. It is straightforward to check that $f$ is a $G$-bundle isomorphism.
Finally, we show that $f^{*}A=A_{\rho}$. For an arbitrary path $\alpha$ on $M$ and $[\delta,g]\in P_{\rho}$ such that $\delta(1)=\alpha(0)$, the horizontal lift of $\alpha$ at $[\delta,g]$ is $\tilde{\alpha}(t)=[\alpha_{(t)}\delta,g]$. Here, $\alpha_{(t)}$ is defined as before, which is the part of $\alpha$ starting at $\alpha(0)$ and ending at $\alpha(t)$. So $f\left([\alpha_{(t)}\delta,g]\right)$ is the ending point of the path on $P$ which is the horizontal lift of $\alpha_{(t)}\delta$ starting at $u_{0}g$. In other words, $f\circ\tilde{\alpha}(t)$ is on the horizontal lift of $\alpha$ starting at $f\left([\delta,g]\right)$ and on the fiber of $\alpha(t)$. Therefore, $f\circ\tilde{\alpha}$ is a horizontal path on $P$. Hence, $f_{*}$ sends horizontal vectors to horizontal vectors and $f^{*}(A)=A_{\rho}$. ∎
@@
Now we turn to the case that $\mathbb{R}/H$ is not a Lie group. In this case $H^{+}$ is non-empty and has no minimal number.
@@
When $H^{+}$ is non-empty and has no minimal number, a cone-flat connection is flat.
@@
By Lemma 4.5, there exists $\xi\in\mathfrak{g}$ such that for any contractible loop $\gamma$ and disk $D$ with $\partial D=\gamma$, we have $hol(\gamma)=-\mathrm{exp}\Big{(}\int_{D}\zeta\cdot\xi\Big{)}$. If $\xi\neq 0$, there exists some small enough $t_{0}$ such that $\mathrm{exp}(t\xi)\neq e$, the identity element of $G$, for any $0<t<t_{0}$.
Since $H^{+}$ has no minimal number, there exists a closed sphere $\Sigma\subset M$ such that $\int_{\Sigma}\zeta=t$ for some $0<t<t_{0}$. Let $\gamma$ be a constant loop at some point $p\in\Sigma$, and $D=\Sigma\setminus\{p\}$. Then $\partial D=\gamma$ so that $hol(\gamma)=-\mathrm{exp}\Big{(}\int_{D}\zeta\cdot\xi\Big{)}$. But $hol(\gamma)=e$ as $\gamma$ is the identity loop. On the other hand, $\int_{D}\zeta=\int_{\Sigma}\zeta=t$. This implies $\mathrm{exp}(t\xi)=e$, which is a contradiction.
Therefore, $\xi$ must be zero and the connection is flat. ∎
@@
We point out that the classification of cone-flat bundles generally depends on the choice of $\zeta$. In Theorem 4.1, the $\zeta$ dependence explicitly appears in the definition of $H$ in (4.3). Below, we will give an example of the classification and its dependence on $\zeta$ in the simple example of $U(1)$ bundles over the four-dimensional torus $M=T^{4}$.
@@
We describe $T^{4}$ as $\mathbb{R}^{4}/\sim$, with the identification $(x_{1},x_{2},x_{3},x_{4})\sim(x_{1}+a,x_{2}+b,x_{3}+c,x_{4}+d),a,b,c,d\in\mathbb{Z}$, and $S^{1}$ as $\{z\in\mathbb{C}\mid|z|=1\}$. Let
@@
be a closed 2-form with $c_{1},c_{2}\in\mathbb{R}\setminus\{0\}$. Then the equivalent classes of $U(1)$ bundles with a cone-flat connection are the conjugacy classes of morphisms $\rho:\Gamma\to S^{1}$, where $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$. Since $S^{1}$ is an abelian group, this classification of cone-flat connections is just the morphisms $\rho:\Gamma\to S^{1}$.
Since $\pi_{2}(T^{4})$ is trivial, $\Gamma$ is an $\mathbb{R}$-extension of $\pi_{1}(T^{4})$. To describe its group structure explicitly, let $a_{i}$ be the straight line path in $\mathbb{R}^{4}$ starting at the origin and ending at the point where the $i$-th coordinate, $x_{i}=1\,$, and $x_{j}=0$ for $j\neq i\,$. When projected to $T^{4}$, $a_{1},\ldots,a_{4}$ become the generators of $\pi_{1}(T^{4})$. Although $\pi_{1}(T^{4})$ is abelian, $a_{1},\ldots,a_{4}$ may no longer commute in its $\mathbb{R}$-extension $\Gamma$. Recall $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is the space of closed paths quotient by contractible closed paths with a special property. For a contractible path $b$, there is an oriented disk $D$ such that $\partial D=b$ and $\partial D$ has the same orientation as $b$. Let $|b|=\int_{D}\zeta$, and we note that $|b|$ is independent of the choice of $D$. In defining $\Gamma$, we take the quotient over contractible paths whose integral, $|b|=\int_{D}\zeta\,=0\,$. Moreover, contractible loops $b$ and $b^{\prime}$ would represent different classes in $\Gamma$ if and only if $|b|\neq|b^{\prime}|$. So, a class of contractible loops $b\in\Gamma$ can be represented by a real number $|b|$.
On $T^{4}$ then, we we can think of $\Gamma$ being generated by $\{a_{1},a_{2},a_{3},a_{4},|b|\}$, with multiplication defined by
@@
Now, the possible morphisms of $\rho:\Gamma\to S^{1}$ is dependent on whether $\frac{c_{1}}{c_{2}}\in\mathbb{Q}$. Since $S^{1}$ is abelian,
@@
When $\frac{c_{1}}{c_{2}}\in\mathbb{Q}$, $\{(pc_{1}+qc_{2})\,|\,p,q\in\mathbb{Z}\}$ has a minimal positive number $c_{0}$. Then $\rho(b)$ must have the form $e^{2\pi i\frac{n|b|}{c_{0}}}$ for some $n\in\mathbb{Z}$. If $\rho(b)=e^{2\pi i\frac{n|b|}{c_{0}}}$, the Euler class of the circle bundle is $\frac{n}{c_{0}}\zeta$, and the definition of $\rho(a_{i})$ determines the connection chosen.
When $\frac{c_{1}}{c_{2}}\notin\mathbb{Q}$, $\{(pc_{1}+qc_{2})\,|\,p,q\in\mathbb{Z}\}$ is dense in $\mathbb{R}$. So $\rho(b)$ must be 1 for any $|b|\in\mathbb{R}$. Thus, the classification is only dependent on $\rho(a_{i})$, and becomes equivalent to the classification of flat connections. Actually, the Euler class in this case is $c\,\zeta$ for some $c\in\mathbb{R}$ but it must be integral. So the only possible $c$ is 0, i.e. every cone-flat connection is flat.
@@
When $M$ is simply-connected, the Hurewicz homomorphism $\pi_{2}(M)\to H_{2}(M,\mathbb{Z})$ is surjective. Assuming $\zeta$ is not $d$-exact, $H$ is discrete if and only if there exists some non-zero constant $c$ such that $c[\zeta]\in H^{2}(M,\mathbb{Z})$. Therefore, if such $c$ exists, then $\Gamma=\mathbb{R}/\mathbb{Z}$ and the classification of isomorphism classes of cone-flat connections on $G$-bundles is given by the conjugacy classes of $\text{Hom}(U(1),G)$. If such $c$ does not exist, then $\Gamma$ is trivial and then there does not exist non-trivial cone-flat connections.
As an application of this observation, let $M$ be simply-connected and closed. If $M$ is also a projective manifold, then we can choose the Kähler form to be integral class and also $\zeta\in H^{2}(M,\mathbb{Z})$. The isomorphism classes of cone-flat connections on $G$-bundles is then given by the conjugacy classes of $\text{Hom}(U(1),G)$. If however $M$ is a non-projective Kähler manifold and we still let $\zeta$ be the Kähler form, then there is no non-zero constant $c$ such that $c[\zeta]\in H^{2}(M,\mathbb{Z})$ (see, for a general reference, [Huybrechts]*Corollary 5.3.3) and the isomorphism classes of cone-flat connections would be trivial.

### v2

@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
For any contractible loop $\gamma\in\Psi^{a}_{a,\zeta}(M)$ and a disk $D$ such that $\partial D=\gamma$, it follows from the definition of $\Psi^{a}_{a,\zeta}(M)$ that $\int_{D}\zeta=0$. By Lemma 4.5, the holonomy along $\gamma$ in $(P,A)$ is given by $hol(\gamma)=\mathrm{exp}\Big{[}\Big{(}\int_{D}\zeta\,\Big{)}\,\xi\Big{]}=e$. Hence, we define the following homomorphism:
$\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G,\quad\gamma\mapsto\rho(\gamma)=hol(\gamma)\,.$ (4.6)
We will show that the resulting $(P_{\rho},A_{\rho})$ is equivalent to $(P,A)$. We start by defining a $G$-bundle isomorphism between the two principal bundles.
Given $[\delta,g]\in P_{\rho}$, let $\tilde{\delta}$ be the horizontal lift of $\delta$ in $P$ with $\tilde{\delta}(0)=u_{0}$. We define
$f:P_{\rho}\to P,\quad[\delta,g]\mapsto\tilde{\delta}(1)\,g\,.$ (4.7)
Note that the definition of $f$ utilizes the horizontal lift determined by the connection $A$ in $P$. Let us show that this map is well-defined. Suppose $[\delta_{1},g_{1}]$ and $[\delta_{2},g_{2}]$ represent the same class in $P_{\rho}$. Then there must exists a $\gamma_{0}\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ such that
$(\delta_{1},g_{1})=(\delta_{2}\gamma_{0},\rho(\gamma_{0}^{-1})g_{2})\sim(\delta_{2},g_{2})\,.$ (4.8)
Thus, $\gamma_{0}^{-1}\delta_{2}^{-1}\delta_{1}\in\Psi^{a}_{a,\zeta}(M)$, and its holonomy in $(P,A)$ is trivial according to Lemma 4.5. Moreover, we have
$hol(\delta_{2}^{-1}\delta_{1})=hol(\gamma_{0})=\rho(\gamma_{0})\,,$ (4.9)
where the last equality follows from our definition of $\rho$ in (4.6). Now let $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ denote the horizontal lift in $(P,A)$ of $\delta_{1}$ and $\delta_{2}$, respectively, starting at $u_{0}$. Since $\tilde{\delta}_{1}$ and the horizontal lift of $\delta_{2}\gamma_{0}=\delta_{2}(\delta_{2}^{-1}\delta_{1})$ starting at $u_{0}$ have the same ending point, so does $\tilde{\delta}_{1}$ and the horizontal lift of $\delta_{2}$ starting at $u_{0}\cdot hol(\gamma_{0})$. Hence, the ending points of $\tilde{\delta}_{1}$ and $\tilde{\delta}_{2}$ satisfy $\tilde{\delta}_{1}(1)=\tilde{\delta}_{2}(1)\cdot hol(\gamma_{0})$. Together with (4.7)-(4.9), we find
@@
Hence, $f$ is well-defined. It is also straightforward to check that $f$ is a $G$-bundle isomorphism.
Finally, we check that the definition of $A_{\rho}$ described earlier in Step 1 is consistent with $f^{*}A=A_{\rho}$. For $[\delta,g]\in P_{\rho}$ and an arbitrary path $\alpha$ on $M$ such that $\alpha(0)=\delta(1)$, the horizontal lift of $\alpha$ at $[\delta,g]$ is $\tilde{\alpha}(t)=[\alpha_{(t)}\delta,g]$. Here again, $\alpha_{(t)}$ is the subpath of $\alpha$ starting at $\alpha(0)$ and ending at $\alpha(t)$. Notice that $f\left([\alpha_{(t)}\delta,g]\right)=(\widetilde{\alpha_{(t)}\delta})(1)\,g$ is the ending point of the path on $P$ that is the horizontal lift of $\alpha_{(t)}$ in $(P,A)$ starting at $f\left([\delta,g]\right)=\tilde{\delta}(1)\,g$. Clearly then, $f\circ\tilde{\alpha}$ is a horizontal path on $(P,A)$. Hence, $f_{*}$ sends horizontal vectors to horizontal vectors, and therefore, $f^{*}(A)=A_{\rho}$. ∎
@@
We now turn to the case when $\mathbb{R}/H$ is not a Lie group. In this case, $H^{+}$ is non-empty and has no minimal number.
@@
When $H^{+}$ is non-empty and has no minimal number, a cone-flat connection is a flat connection.
@@
By Lemma 4.5, there exists some $\xi\in\mathfrak{g}$ such that for any contractible loop $\gamma$ and disk $D$ with $\partial D=\gamma$, we have $hol(\gamma)=\mathrm{exp}\Big{[}\Big{(}\int_{D}\zeta\,\Big{)}\,\xi\Big{]}$.
Let us show that $\xi$ must be zero. For if $\xi\neq 0$, then there exists some small enough $t_{0}$ such that for any $0<t<t_{0}\,$, $\mathrm{exp}(t\,\xi)\neq e\,$ the identity element of $G$. Now since $H^{+}$ has no minimal number, there exists a closed sphere $S\subset M$ such that $\int_{S}\zeta=t$ for some $0<t<t_{0}$. Let $\gamma$ be a constant loop at some point $p\in S$, and $D=S\setminus\!\{p\}\,$. Then $\partial D=\gamma$ so that $hol(\gamma)=\mathrm{exp}\Big{[}\Big{(}\int_{D}\zeta\,\Big{)}\,\xi\Big{]}$. But $hol(\gamma)=e$ as $\gamma$ is the identity loop. On the other hand, $\int_{D}\zeta=\int_{S}\zeta=t\,$, and therefore, $\mathrm{exp}(t\,\xi)=e$, which gives a contradiction. Thus, we conclude that $\xi$ is zero.
With $\xi=0$, the holonomy of any contractible loop is trivial. Hence, the curvature vanishes and the connection is flat. ∎
@@
We point out that the classification of cone-flat bundles generally depends on the choice of $\zeta$. In Theorem 4.1, the $\zeta$ dependence explicitly appears in the definition of $H$ in (4.3). Below, we will work out the classification and demonstrate its dependence on $\zeta$ in the simple example of $U(1)$ bundles over the four-dimensional torus $M=T^{4}$.
@@
We describe $T^{4}$ as $\mathbb{R}^{4}/\sim$, with the identification $(x_{1},x_{2},x_{3},x_{4})\sim(x_{1}+a,x_{2}+b,x_{3}+c,x_{4}+d)$ where $a,b,c,d\in\mathbb{Z}$, and $U(1)$ as $\{z\in\mathbb{C}\mid|z|=1\}$. Let
@@
be a closed 2-form with $c_{1},c_{2}\in\mathbb{R}\setminus\{0\}$. By Theorem 4.1, the equivalent classes of $U(1)$ bundles with a cone-flat connection are in 1-1 correspondence with homomorphisms $\rho:\Gamma\to U(1)$, where $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$.
Since $\pi_{2}(T^{4})$ is trivial, $\Gamma$ is an $\mathbb{R}$-extension of $\pi_{1}(T^{4})$. To describe its group structure explicitly, let $a_{i}$ for $i=1,2,3,4$ be the straight line path in $\mathbb{R}^{4}$ starting at the origin and ending at the point where the $i$-th coordinate, $x_{i}=1\,$, and $x_{j}=0$ for $j\neq i\,$. When projected to $T^{4}$, $\{a_{1},a_{2},a_{3},a_{4}\}$ become the generators of $\pi_{1}(T^{4})$. Although $\pi_{1}(T^{4})$ is abelian, the elements of $\{a_{1},a_{2},a_{3},a_{4}\}$ may no longer commute in $\Gamma$.
For a contractible loop $b$, there is an oriented disk $D$ such that $\partial D=b$ and $\partial D$ has the same orientation as $b$. Let $|b|=\int_{D}\zeta$, and we note that $|b|$ is independent of the choice of $D$. Now recall that $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\,$ is a quotient of loops by contractible ones whose integral, $|b|=\int_{D}\zeta\,=0\,$. Moreover, contractible loops $b$ and $b^{\prime}$ would represent different classes in $\Gamma$ if and only if $|b|\neq|b^{\prime}|$. So, a class of contractible loops $b\in\Gamma$ can be represented by the real number $|b|$.
On $T^{4}$ then, we can think of $\Gamma$ as being generated by $\{a_{1},a_{2},a_{3},a_{4},|b|\}$, with multiplication defined by
@@
Now, the possible homomorphisms of $\rho:\Gamma\to U(1)$ is dependent on whether $\frac{c_{1}}{c_{2}}\in\mathbb{Q}$. Since $U(1)$ is abelian,
@@
When $\frac{c_{1}}{c_{2}}\in\mathbb{Q}$, $\{(pc_{1}+qc_{2})\,|\,p,q\in\mathbb{Z}\}$ has a minimal positive number $c_{0}$. Then $\rho(b)$ must have the form $e^{2\pi i\frac{n|b|}{c_{0}}}$ for some $n\in\mathbb{Z}$. In this case, the Euler class of the circle bundle is $\frac{n}{c_{0}}\zeta$, and the choice of $\rho(a_{i})$ for $i=1,\dots,4$ determines the connection.
When $\frac{c_{1}}{c_{2}}\notin\mathbb{Q}$, $\{(pc_{1}+qc_{2})\,|\,p,q\in\mathbb{Z}\}$ is dense in $\mathbb{R}$. So $\rho(b)$ must be 1 for any $|b|\in\mathbb{R}$. Thus, the classification is only dependent on $\rho(a_{i})$, and becomes equivalent to the classification of flat connections. Actually, the Euler class in this case is $c\,\zeta$ for some $c\in\mathbb{R}$. But the Euler class is an integral class, and hence, the only possible $c$ is 0, i.e. every cone-flat connection is flat.
@@
When $M$ is simply-connected, the Hurewicz homomorphism $\pi_{2}(M)\to H_{2}(M,\mathbb{Z})$ is surjective. Assuming $\zeta$ is not $d$-exact, $H$ is discrete if and only if there exists some non-zero constant $c$ such that $c[\zeta]\in H^{2}(M,\mathbb{Z})$. Therefore, if such $c$ exists, then $\Gamma=\mathbb{R}/\mathbb{Z}$ and the classification of isomorphism classes of cone-flat connections on $G$-bundles is given by the conjugacy classes of $\text{Hom}(S^{1},G)$. If such $c$ does not exist, then $\Gamma$ is trivial and then there does not exist non-trivial cone-flat connections.
As an application of this observation, let $M$ be simply-connected and closed. If $M$ is also a projective manifold, then we can choose $\zeta$ to be the Kähler form which is an integral class, and so, $\zeta\in H^{2}(M,\mathbb{Z})$. The isomorphism classes of cone-flat connections on $G$-bundles is then given by the conjugacy classes of $\text{Hom}(S^{1},G)$. If however $M$ is a non-projective Kähler manifold and we still let $\zeta$ be the Kähler form, then there is no non-zero constant $c$ such that $c[\zeta]\in H^{2}(M,\mathbb{Z})$ (see, for a reference, [Huybrechts]*Corollary 5.3.3) and the isomorphism classes of cone-flat connections would be trivial.